OpenStudy (anonymous):

Solve this inequality and write the answer in interval notation: (4)/(x-3)-(1)/(2x)≥0

OpenStudy (anonymous):

Do what you said before.. Put them over a common denominator.

OpenStudy (dumbcow):

combine fractions common denominator is (x-3)(2x) 8x-(x-3) / 2x(x-3) >=0 simplifies to 7x+3 >= 0 7x >= -3 x >= -3/7

OpenStudy (anonymous):

thank you. at least someone knows how to show their work.

OpenStudy (anonymous):

Nobody cares if WE know how to do it. The question is do YOU know how to do it. If not then you will have to learn. And you won't do that by simply following along someone else's thought process.

OpenStudy (anonymous):

i think i will lol

OpenStudy (anonymous):

no im kidding. your right

OpenStudy (anonymous):

Really? Cause dumbcow's answer is wrong. And if you write it down then you are also wrong.

OpenStudy (anonymous):

if his answer is wrong, then what is the right answer?

OpenStudy (dumbcow):

very true... repeat these steps on your own to practice

OpenStudy (anonymous):

What we said the last time you asked this.. \([-3/7, 0) \bigcup (3,\infty)\)

OpenStudy (anonymous):

And you wanted us to show work. But I would rather walk you through the process of finding the answer yourself. I think it will be more instructive.

OpenStudy (anonymous):

that answer is different than the answer you gave me last time, so how do i know that your answer is right?

OpenStudy (anonymous):

I didn't give you an answer last time. Someone else did. I didn't check it because I figured you wanted to work it out anyway.

OpenStudy (anonymous):

Do you want to work it out? This is why it doesn't help for someone else to show you their work. You have to work it out for yourself or you cannot be sure that the answer is correct.

OpenStudy (dumbcow):

no polpak is correct, i didn't look at conditions for denominator being 0

OpenStudy (anonymous):

ok. so can you show me the work on how to do it right

OpenStudy (anonymous):

Certainly.

OpenStudy (anonymous):

Start again with: \[\frac{4}{x-3} - \frac{1}{2x} \ge 0\] And put the fractions over a common denominator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What do you have?